OpenStudy (anonymous):

number of primes less than 30

OpenStudy (anonymous):

2, 3, 5, 7, 11, 13, 17, 19, 23, 29

OpenStudy (anonymous):

1 is not a prime number

OpenStudy (anonymous):

2 to the power of 4 minus 5

OpenStudy (anonymous):

another one

OpenStudy (anonymous):

11

OpenStudy (anonymous):

yay i got that right!!! (there are more im checkin if they r right

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

9.4+ 8 3/5

OpenStudy (radar):

9 4/10 + 8 3/5 9 4/10 + 8 6/10 17 + 10/10=17 + 1 = 18